If the width of the rectangle is 3 more than the width. What are the dimensions of the rectrangle if its perimeter is 23?
Answer by macston(5194)

Problem not possible as written: width cannot be 3 more than the width.
If length is 3 more than the width:
W=width; L=length=3W; P=perimeter=23
P=2(L+W) Substitute for L.
23=2(W+3+W) Divide each side by 2.
11.5=2W+3 Subtract 3 from each side
8.5=2W Divide each side by 2
4.25=W ANSWER 1: The width is 4.25 units.
L=W+3=4.25+3=7.25 ANSWER 2: The length is 7.25 units.
CHECK:
P=2(L+W)
23=2(7.25+4.25)
23=2(11.5)
23=23